# Collaborative Multi-Agent Optimization for Personalized Memory System

### 13 Mar 2026 v1

@@ 4.3 Adaptive Credit Assignment Mechanism
To align local improvements with global performance during joint optimization, we propose an adaptive credit assignment mechanism to allocate global system rewards to local agents. Instead of assigning global rewards equally to each agent (Hong et al., 2025; Zhao et al., 2026), CoMAM quantifies agents’ varying contributions to the system and treats them as adaptive weights for allocation.
Formally, for each input $(\mathcal{H},q)$, we sample a group of MDP trajectories $\{\tau_{i}\}_{i=1}^{G}$, collecting the local rewards for each agent $n$, $\mathcal{R}_{\text{local},n}=\{r_{n}^{(i)}\}_{i=1}^{G}$, and the corresponding global rewards, $\mathcal{R}_{\text{global}}=\{r_{3}^{(i)}\}_{i=1}^{G}$, where $n\in\{0,1,2\}$.
To quantify agent $n$’s contribution to the memory system, we compute the ranking consistency score between $\mathcal{R}_{\text{local},n}$ and $\mathcal{R}_{\text{global}}$ using the Normalized Discounted Cumulative Gain (NDCG) metric. Specifically, the NDCG-based consistency score for agent $n$ is defined as:
@@
where the global reward ranking $\sigma(\mathcal{R}_{\text{global}})$ serves as the ideal ranking for the local reward ranking $\sigma(\mathcal{R}_{\text{local},n})$. We then normalize $v_{n}$ to obtain adaptive credit weights $w_{n}$ that satisfy $\sum_{n}w_{n}=1$:
$w_{n}=\frac{\exp(v_{n})}{\sum_{n^{\prime}}\exp(v_{n^{\prime}})}.$ (14)
The final optimization signal for agent $n$ is the combination of its local rewards and allocated global credit, weighted by the adaptive weight $w_{n}$:
@@
After the adaptive credit assignment, we apply the GRPO algorithm (as defined in Sec 3.2) on each agent based on the integrated rewards $r_{\text{final}}$. Formally, the objective function for the agent $n$ is defined as:
$\mathcal{J}_{GRPO}\left(\pi_{\theta}^{n}\right)=$ (16)
$\frac{1}{G}\sum_{i=1}^{G}\min\left[\rho_{n}^{(i)}A_{n}^{(i)},\;\text{clip}\left(\rho_{n}^{(i)},\;1\pm\epsilon\right)A_{n}^{(i)}\right],$
where $\rho_{n}^{(i)}=\frac{\pi_{\theta}^{n}\left(a_{n}^{(i)}\mid s_{n}^{(i)}\right)}{\pi_{\theta_{\text{old}}}^{n}\left(a_{n}^{(i)}\mid s_{n}^{(i)}\right)}$ is the importance sampling ratio, the advantage $A_{n}^{(i)}$ for each agent’s action $a_{n}^{(i)}$ is computed by normalizing its reward $r_{\text{final},n}^{(i)}$. By optimizing each agent with both local task-specific rewards and allocated global system rewards, we can align local improvements with global performance and foster collaboration.
@@ 5 Experiments
  • RQ1: How does CoMAM perform against leading baselines of memory systems for personalization?
@@
  • RQ3: What is the impact of regularizing agents’ execution as an MDP for jointly optimizing the memory system?
  • RQ4: What is the contribution of adaptive credit assignment for facilitating collaboration in the memory system?
@@
Table 1: Query-Answering Accuracy on the PersonaMem Benchmark across 32K, 128K, and 1M context length settings. Bold indicates the best performance, and underlining denotes the second best.
@@ 5.1 Experimental Settings
We conduct experiments on the PersonaMem benchmark (Jiang et al., 2025), a large-scale dataset including over 180 long-term user-LLM interaction histories and about 6,000 multiple-choice questions as subsequent queries. The benchmark is split into three experimental settings with varying history lengths (i.e., 32K, 128K, and 1M tokens), each paired with subsequent queries spanning seven distinct types: (1) Recall user-shared facts, (2) Suggest new ideas, (3) Acknowledge latest user preferences, (4) Track full preference evolution, (5) Revisit reasons for preference updates, (6) Provide preference-aligned recommendations, and (7) Generalize to new scenarios. Compared to existing benchmarks (Maharana et al., 2024; Wu et al., 2025), PersonaMem’s strengths in long context history and personalized query type enable rigorous assessment of memory systems in personalized conversational scenarios. Dataset processing details are provided in Appendix B.1.
@@
We conduct a comprehensive evaluation of CoMAM against multiple leading baselines, including methods without a memory system “No memory”, methods of building memory systems via dedicated prompts “Prompt-based memory”, and methods of optimizing the memory systems with RL “RL-based memory”. Detailed introductions and implementations of baselines are provided in Appendix B.
  • No memory: “Base” represents directly injecting raw conversations as LLMs’ input for query answering; RAG (Lewis et al., 2020) answers queries by retrieving raw conversational history;
  • Prompt-based memory: A-Mem (Xu et al., 2025) instantiates a memory system with construction, evolution, and retrieval agents; CAM (Li et al., 2025) and MemoryBank (Zhong et al., 2024) build multi-level (fine-to-coarse granularity) memory via multi-agent dedicated prompts.
  • RL-based memory: Mem1 (Zhou et al., 2026) uses RL to optimize agents for memory searching during reasoning; Memory-R1 (Yan et al., 2025b) separately optimizes memory construction and retriever agents via RL.
@@
To evaluate CoMAM’s effectiveness and generalizability, we conduct experiments on PersonaMem across its three history length settings (i.e., 32K, 128K, and 1M tokens) using LLM backbones from the Qwen (Bai et al., 2023) and Llama (Touvron et al., 2023) families. Specifically, the extraction agent, profile agent, and reward model $\mathcal{V}$ are instantiated with Qwen2.5-3B-Instruct and Llama-3.2-3B-Instruct, while the retrieval agent leverages Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct, with model sizes tailored to each agent’s task complexity. All LLM backbones are configured with a 32K-token context window, and we adopt the average query-answer accuracy as the personalization evaluation metric. Table 1 presents the main results for both Qwen and Llama; other tables show only Qwen for simplicity. All tabulated results are averages of at least three independent runs with distinct random seeds. More details are provided in the Appendix B.3.
@@ 5.2 Overall Performance (RQ1)
To validate CoMAM’s efficacy for personalized LLMs, we compare its query-answer accuracy against leading memory system baselines on the PersonaMem benchmark. As shown in Table 1, CoMAM delivers superior performance across all context length configurations (32K, 128K, and 1M) and two LLM backbone families (Qwen and Llama), fully confirming CoMAM’s effectiveness and generalizability in personalized conversational scenarios. Notably, RL-based memory baselines outperform prompt-based memory systems, validating that reinforcement learning is effective for optimizing memory system agents. For instance, the top prompt-based baseline CAM (Li et al., 2025) achieves only 0.450.45-$0.53$ accuracy across contexts, while the RL-based Memory-R1 (Yan et al., 2025b) maintains a higher accuracy range of 0.570.570.610.61. Moreover, CoMAM outperforms the state-of-the-art baseline Memory-R1, which employs independent agent optimization. This highlights the critical importance of joint optimization for multi-agent memory systems to achieve collaboration. Figure 4 details Qwen-based results for the seven question types, and we observe that CoMAM performs well in both detailed factual recall (e.g., Type 1) and abstract preference-based question types (e.g., Type 6), providing additional validation of its efficacy.
@@
Table 3: Ablation study of MDP regularization for heterogeneous, asynchronous agents. The “no train” variant denotes a memory system composed of agents merely guided by prompting.
@@ 5.3 Ablation for Agents (RQ2)
To verify the necessity and contribution of the three core agents (Extraction Agent, Profile Agent, and Retrieval Agent), we designed targeted ablation experiments, with results presented in Table 2. First, the “Base” baseline (without memory agents) achieves the lowest performance across all context lengths, confirming the significance of memory systems for long-context personalized conversations. Second, removing any agent can lead to significant performance degradation, whether the Extraction Agent (“w/o Ex”) that builds factual fine-grained memories, the Profile Agent (“w/o Pr”) that captures core user preferences, or the Retrieval Agent (“w/o Re”) that searches for memories to answer queries. These results demonstrate that all three agents play complementary and unique roles in supporting the personalized memory system. Furthermore, the “w/o train” variants (full agent architecture with one agent left untrained) provide additional validation: untraining any single agent results in 2–8% performance drops compared to CoMAM, highlighting that optimizing these agents is critical to building an effective memory system.
5.4 Ablation for MDP Regularization (RQ3)
To address heterogeneity and asynchrony for joint optimization, we regularize agents’ execution as a sequential MDP trajectory. To validate MDP Regularization’s efficacy, we design “single policy” (all agents share one policy across tasks (Yu et al., 2026; Bo et al., 2024b)) and “heterogeneous policy” (each agent has a dedicated policy per task) variants, with two training strategies: “independent RL” (fine-tunes one agent’s policy at a time, others fixed without updates) and “MDP RL” (simultaneously updates policies via MDP trajectory for joint optimization). As shown in Table 3, “single policy” variants consistently underperform “heterogeneous policy” variants, confirming that a shared policy undermines agent specialization by integrating diverse task-specific objectives. Furthermore, “+MDP RL” outperforms “independent RL” across both policy paradigms, validating that MDP regularization explicitly models inter-agent sequential dependencies and enables both local task specialization and global collaboration.
@@ 5.5 Ablation for Adaptive Credit Assignment (RQ4)
To address RQ4, we design diverse credit assignment strategies to validate CoMAM’s adaptive mechanism. As shown in Table 4, optimizing agents solely with local task-specific rewards (“MDP RL w/ L”) or global rewards (“MDP RL w/ G”) yields performance improvements over the “no train” variant and “independent RL” variant, but lags behind variants that integrate global rewards with local rewards (“MDP RL w/ LG” and “Ours”). This demonstrates the necessity of combining local task rewards and global rewards to balance agent specialization on local tasks and collaboration for global performance. Notably, “Ours” (which is equipped with adaptive credit assignment) outperforms “MDP RL w/ LG” (which allocates equal shares of global rewards to local tasks), confirming the critical value of adaptively aligning local improvements with global performance. As illustrated in Figure 5, our adaptive weight $w_{n}$ for credit assignment outperforms any fixed weights, further validating the efficacy of the adaptive mechanism. Moreover, adaptive credit assignment can also enhance agents’ task-specific ability: our method outperforms both the “w/ LG” and “w/ L” on local task performance, validating this effect.
@@
5.6 Efficiency (RQ5)
We further evaluate the training efficiency of CoMAM by comparing the convergence steps of joint optimization and independent optimization across 32K, 128K, and 1M context length settings, with results presented in Table 5. The total convergence steps serve as the core metric to reflect training efficiency, where fewer steps indicate higher efficiency in reaching stable performance. Specifically, independent optimization trains each agent separately, with total steps being the sum of individual convergence steps for the Extraction, Profile, and Retrieval Agents. In contrast, joint optimization trains all agents in parallel, with total steps determined by the slowest-converging agent, resulting in fewer total steps. While individual agents under joint training may take more steps (e.g., 58 vs. 48 steps for the Extraction Agent at the 32K setting), parallel training shortens the total training duration. Notably, inference latency is decoupled from the training paradigm: agents optimized by joint or independent optimization achieve comparable inference efficiency at deployment. Yet joint optimization delivers superior global performance via inter-agent collaboration, highlighting the distinct advantage of joint multi-agent optimization for personalized memory systems.
@@ 6 Conclusion
In this paper, we propose CoMAM, a collaborative reinforcement learning framework for jointly optimizing agents in personalized memory systems. It tackles the issue that independent agent optimization can not ensure global system performance without cross-agent collaboration. To enable joint optimization, we regularize the asynchronous executions of heterogeneous agents as a sequential MDP, which embeds inter-agent dependencies into state transitions and yields both local and global rewards. To align local improvement with global performance, we introduce an adaptive credit assignment mechanism that allocates global rewards to individual agents based on group-level ranking consistency between global and local rewards. Experiments on long-term personalized conversational benchmarks demonstrate that CoMAM outperforms leading memory baselines, validating the effectiveness of our collaborative RL framework. The limitation is discussed in Appendix E.
Impact Statements
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.

### 27 Apr 2026 v2

@@ 4.3 Adaptive Credit Assignment Mechanism
While modeling inter-agent dependencies during joint training enables co-adaptation, relying solely on a shared global reward provides limited and ambiguous guidance for optimizing specialized agents hong2025multi ; naive_assign . In particular, when agents cannot distinguish their individual impact on the overall outcome, they receive indistinguishable learning signals, which prevent the development of effective role-specific policies and consequently hinder agents’ collaboration in the memory system. To address this limitation, we introduce an adaptive credit assignment mechanism that distributes the global reward among agents in proportion to their estimated contribution, providing more informative training signals for each specialized agent.
Formally, for each input $(\mathcal{H},q)$, we sample a group of MDP trajectories $\{\tau_{i}\}_{i=1}^{G}$, and collect local rewards for each agent $n$, $\mathcal{R}_{\text{local},n}=\{r_{n}^{(i)}\}_{i=1}^{G}$, along with the corresponding global rewards, $\mathcal{R}_{\text{global}}=\{r_{3}^{(i)}\}_{i=1}^{G}$, where $n\in\{0,1,2\}$. To estimate each agent’s contribution to the global system, we measure the alignment between its local and global rewards using a ranking-based consistency metric, instantiated as Normalized Discounted Cumulative Gain (NDCG) NDCG_1 :
@@
where $\sigma(\cdot)$ denotes the ranking operator. Intuitively, this metric captures whether trajectories with higher global rewards are also associated with higher local rewards for a given agent. Agents whose local reward rankings are more consistent with global outcome rankings receive higher credit, as their behaviors are more aligned with improving overall performance. Moreover, compared to value-based metrics, ranking-based measures are more robust under sparse and noisy rewards, as they rely on relative ordering rather than absolute magnitudes ranking_reward . The resulting scores are normalized to obtain adaptive weights $w_{n}$ with $\sum_{n}w_{n}=1$: $w_{n}=\frac{\exp(v_{n})}{\sum_{n^{\prime}}\exp(v_{n^{\prime}})}$. Each agent’s final optimization signal combines its local reward with its adaptively weighted share of the global reward:
@@
We then optimize each agent simultaneously using the GRPO algorithm (as introduced in Sec. 3.2) on these integrated rewards: $\mathcal{J}_{\text{GRPO}}\left(\pi_{\theta}^{n}\right)=\frac{1}{G}\sum_{i=1}^{G}\min\Big[\rho_{n}^{(i)}A_{n}^{(i)},\;\text{clip}(\rho_{n}^{(i)},1\pm\epsilon)A_{n}^{(i)}\Big]$, where $\rho_{n}^{(i)}$ is the importance sampling ratio and $A_{n}^{(i)}$ the normalized advantage computed from $r_{\text{final},n}^{(i)}$. By leveraging both local task-specific rewards and adaptively weighted global rewards, this mechanism fosters collaboration while preserving agent specialization.
@@ 5 Experiments
  • RQ1: How does CoMAM perform against leading memory systems?
@@
  • RQ3: What is the impact of inter-agent dependency modeling for joint optimization?
  • RQ4: What is the contribution of adaptive credit assignment for the specialization agents?
@@
Table 1: Query-Answering Accuracy on the PersonaMem Benchmark across 32K, 128K, and 1M context lengths. Bold indicates the best performance, and underlining denotes the second best.
@@
Table 2: Results on LongMemEval using Qwen2.5-7B.
Method $F_{1}\uparrow$ $B_{1}\uparrow$ $J_{1}\uparrow$
RAG 18.27 14.57 22.20
SeCom 19.03 16.57 10.00
A-Mem 41.55 36.58 54.80
Mem0 38.44 34.53 46.80
GAM 51.75 48.96 52.00
Ours 62.17 51.88 57.60
@@ 5.1 Experimental Settings
We conduct experiments on the PersonaMem personamem and LongMemEval longmemeval benchmarks. PersonaMem provides long-horizon conversational histories with varying context lengths (i.e., 32K, 128K, and 1M tokens), each paired with seven downstream query types. LongMemEval focuses on multi-session conversational settings, requiring tracking information and performing temporal reasoning across sessions. They offer a more realistic evaluation of long-horizon memory construction and retrieval. Dataset processing details are provided in Appendix B.1.
@@
We conduct a comprehensive evaluation of CoMAM against multiple leading baselines, including methods without a memory system “No memory”, methods of building memory systems via dedicated prompts “Prompt-based memory”, and methods of optimizing the memory systems with RL “RL-based memory”. Detailed information of baselines are provided in Appendix B.
  • No memory: “Base” represents directly injecting raw conversations as LLMs’ input for query answering; RAG RAG answers queries by retrieving raw conversational history;
  • Prompt-based memory: A-Mem xu2025amem instantiates a memory system with construction, evolution, and retrieval agents; CAM li2025cam and MemoryBank memorybank build multi-level (fine-to-coarse granularity) memory via multi-agent dedicated prompts.
  • RL-based memory: Mem1 mem1 uses RL to optimize agents for memory searching during reasoning; Memory-R1 mem-r1 separately optimizes memory construction and retriever agents via RL.
@@
To evaluate CoMAM’s effectiveness and generalizability, we conduct experiments on PersonaMem and LongMemEval using LLM backbones from the Qwen qwen and Llama Llama families. Specifically, the extraction agent, profile agent, and reward model $\mathcal{V}$ are instantiated with Qwen2.5-3B-Instruct and Llama-3.2-3B-Instruct, while the retrieval agent leverages Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct, with model sizes tailored to each agent’s task complexity. All LLM backbones are configured with a 32K-token context window. Following existing work personamem ; mem-r1 ; mem1 , we evaluate memory systems using query-answer accuracy. Specifically, we adopt multiple-choice accuracy for PersonaMem, and $F_{1}$, $B_{1}$, and $J_{1}$ (LLM-as-a-Judge) for LongMemEval. Table 1 presents the main results for both Qwen and Llama; other tables show only Qwen for simplicity. All tabulated results are averages of at least three independent runs with distinct random seeds. More details are provided in the Appendix B.3.
@@ 5.2 Overall Performance (RQ1)
To validate CoMAM’s efficacy for the memory system, we compare its query-answer accuracy against leading memory system baselines on the PersonaMem benchmark across its three history length settings (i.e., 32K, 128K, and 1M tokens). As shown in Table 1, CoMAM delivers superior performance across all context length configurations (32K, 128K, and 1M) and two LLM backbone families (Qwen and Llama), fully confirming CoMAM’s effectiveness and generalizability in long-horizon conversational scenarios. Notably, RL-based memory baselines outperform prompt-based memory systems, validating that reinforcement learning is effective for optimizing memory system agents. For instance, the top prompt-based baseline CAM li2025cam achieves only 0.450.45-$0.53$ accuracy across contexts, while the RL-based Memory-R1 mem-r1 maintains a higher accuracy range of 0.570.570.610.61. Moreover, CoMAM outperforms the state-of-the-art baseline Memory-R1, which employs independent agent optimization. This highlights the critical importance of joint optimization for multi-agent memory systems to achieve collaboration. Figure 6 in Appendix C.2 details Qwen-based results for the seven question types of PersonaMem. We further evaluate CoMAM on the LongMemEval benchmark. As shown in Table 2, our method achieves the best performance across all metrics, consistently outperforming strong baselines such as GAM GAM , further demonstrating its effectiveness in long-context memory retrieval and reasoning scenarios.
@@
Table 4: Ablation study of joint optimization for heterogeneous agents.
@@ 5.3 Ablation for Agents (RQ2)
To verify the necessity and contribution of the three core agents (Extraction Agent, Profile Agent, and Retrieval Agent), we designed targeted ablation experiments, with results presented in Table 4. First, the “Base” baseline (without memory agents) achieves the lowest performance across all context lengths, confirming the significance of memory systems for long-horizon conversations. Second, removing any agent can lead to significant performance degradation, whether the Extraction Agent (“w/o Ex”) that builds factual fine-grained memories, the Profile Agent (“w/o Pr”) that captures core user preferences, or the Retrieval Agent (“w/o Re”) that searches for memories to answer queries. These results demonstrate that all three agents play complementary and unique roles in supporting the memory system. Furthermore, the “w/o train” variants (full agent architecture with one agent left untrained) provide additional validation: untraining any single agent results in 2–8% performance drops compared to CoMAM, highlighting that optimizing these agents is critical to building an effective memory system. We further analyze the sensitivity of the reward coefficients $\alpha$ and $\beta$ in Figure 3, showing that the chosen settings ($\alpha=0.8$, $\beta=0.2$) enable the extraction agent to better preserve key information and the retrieval agent to effectively filter noise.
Figure 3: Sensitivity of CoMAM to the hyperparameters $\alpha$ and $\beta$ as introduced in Sec. 4.1.
5.4 Ablation for Dependency Modeling (RQ3)
To model inter-agent dependencies for joint optimization, we formulate the heterogeneous multi-agent pipeline as a sequential MDP, which enables end-to-end RL. To validate its efficacy, we design “single policy” (all agents share one policy across tasks memagent ; share_one_policy ) and “heterogeneous policy” (each agent has a dedicated policy per task) variants, with two training strategies: “independent RL” (fine-tunes one agent’s policy at a time, others fixed without updates) and “end-to-end RL” (simultaneously updates policies via MDP trajectory for joint optimization). As shown in Table 4, “single policy” variants consistently underperform “heterogeneous policy” variants, confirming that a shared policy undermines agent specialization by integrating diverse task-specific objectives. Furthermore, “+end-to-end RL” outperforms “independent RL” across both policy paradigms. This suggests that sequential MDP modeling during end-to-end RL training acts as a structured abstraction to capture how upstream decisions influence downstream behavior, thereby enabling more effective co-adaptation among agents.
@@ 5.5 Ablation for Adaptive Credit Assignment (RQ4)
To address RQ4, we evaluate different credit assignment strategies to assess the effectiveness of CoMAM’s adaptive mechanism. As shown in Table 6, optimizing agents using only local rewards (“end-to-end RL w/ L”) or only global rewards (“end-to-end RL w/ G”) improves over the “no train” and “independent RL” variants, but underperforms approaches that combine both signals (“end-to-end RL w/ LG” and “Ours”). This suggests that integrating local and global rewards is essential for balancing agent specialization and collaboration. Notably, “Ours”, with adaptive credit assignment, consistently outperforms “end-to-end RL w/ LG”, which uses uniform global rewards, highlighting the advantage of assigning credit based on agents’ contributions. As further shown in Figure 4, adaptive weights $w_{n}$ outperform fixed weighting schemes. Moreover, adaptive credit assignment improves task-specific performance: our method consistently achieves higher local task performance across all agents compared to both “w/ LG” and “w/ L”, indicating stronger role-specific capability. Additional experimental results and analyses of the reward curves are provided in Appendix C.1.
5.6 Efficiency (RQ5)
We further evaluate the training efficiency of CoMAM by comparing the convergence steps of joint optimization and independent optimization across 32K, 128K, and 1M context length settings, with results presented in Table 6. The total convergence steps serve as the core metric to reflect training efficiency, where fewer steps indicate higher efficiency in reaching stable performance. Specifically, independent optimization trains each agent separately, with total steps being the sum of individual convergence steps for the Extraction, Profile, and Retrieval Agents. In contrast, joint optimization trains all agents in parallel, with total steps determined by the slowest-converging agent, resulting in fewer total steps. While individual agents under joint training may take more steps for convergence (e.g., 58 vs. 48 steps for the Extraction Agent at the 32K setting), joint training shortens the total training duration. Importantly, per-step cost is similar in both settings, since frozen agents in independent optimization still participate in the forward pass to complete the pipeline and compute rewards. Thus, the efficiency gain mainly comes from reducing redundant forward passes, as agents update simultaneously in joint optimization. Moreover, inference latency is decoupled from the training paradigm: agents optimized by joint or independent optimization achieve comparable inference efficiency at deployment. Yet joint optimization delivers superior global performance via inter-agent collaboration and adaptive credit assignment, highlighting the distinct advantage of joint multi-agent optimization for memory systems.
@@ 6 Conclusion
In this work, we study the limitations of independently optimizing agents in multi-agent memory systems, highlighting the lack of co-adaptation and the ambiguity of credit assignment under a shared global objective. To address these challenges, we propose CoMAM, a joint optimization framework that enables agents’ co-adaptation via end-to-end reinforcement learning, together with an adaptive credit assignment mechanism for more informative training signals. By modeling the inter-agent dependencies as an MDP and integrating local and global rewards adaptively, CoMAM facilitates collaboration across agents. Experimental results demonstrate consistent improvements over strong baselines, validating the effectiveness of joint optimization with adaptive credit assignment for multi-agent memory systems. The limitation is discussed in Appendix E.